# Protecting Human Activity Signatures in Compressed IEEE 802.11 CSI Feedback

### v1

@@ Abstract
Explicit channel state information (CSI) feedback in IEEE 802.11 conveys transmit beamforming directions by reporting quantized Givens rotation and phase angles that parametrize the right-singular subspace of the channel matrix. Because these angles encode fine-grained spatial signatures of the propagation environment, recent work have shown that plaintext CSI feedback can inadvertently reveal user activity, identity, and location to passive eavesdroppers. In this work, we introduce a standards-compatible differentially private (DP) quantization mechanism that replaces deterministic angular quantization with an $\varepsilon$-DP stochastic quantizer applied directly to the Givens parameters of the transmit beamforming matrix. The mechanism preserves the 802.11 feedback structure, admits closed-form sensitivity bounds for the angular representation, and enables principled privacy calibration. Numerical simulations demonstrate strong privacy guarantees with minimal degradation in beamforming performance.
@@ I Introduction
Beamforming is a core capability of modern multi-antenna wireless systems [goldsmith2005wireless, geraci2025wi], enabling transmitters to dynamically shape and steer radiated energy toward intended receivers. By coherently combining signals across multiple antennas, the transmitter forms highly directional beams that substantially improve the received signal-to-noise ratio (SNR) while suppressing unintended interference—capabilities unattainable with omnidirectional transmission. Introduced into Wi-Fi through the IEEE 802.11n “High Throughput” amendment [ieee80211n2012], beamforming has since become standard in commodity WLAN devices, where the access point (beamformer) computes a channel-dependent steering matrix that optimizes reception quality at the client (beamformee).
Two primary approaches are used to compute this steering matrix [perahia2013next]. In implicit beamforming, the transmitter estimates the downlink channel from uplink pilot transmissions under an assumption of channel reciprocity. While attractive for its simplicity, this approach is often suboptimal in practice due to hardware asymmetries between the transmit and receive chains. By contrast, explicit beamforming—now the mandated method in modern Wi-Fi standards—achieves higher accuracy by relying on channel state information (CSI) fed back by the client. The beamformee estimates the downlink channel, compresses it, and reports the result to the beamformer through dedicated sounding and feedback frames, typically using quantized matrices or precoding matrix indicators (PMIs).
@@ I-A Privacy Concerns in Explicit Beamforming
The privacy risks associated with channel measurements have been extensively demonstrated in the literature on device-free sensing and Wi-Fi-based activity recognition [abdelnasser2015wigest, wang2017device, jiang2021wifi, shenoy2022rf, zhu2025csi]. These works show that subtle CSI variations can reveal human presence, gestures, and even respiration patterns. However, most existing countermeasures—such as temporal smoothing, random antenna selection, or coarse quantization—offer only heuristic protection, lack formal privacy guarantees, and often degrade beamforming performance.
More recently, empirical studies have revealed that standard Wi-Fi beamforming feedback itself can directly leak sensitive information. Liu et al. [liu2024lendmeyourbeam] showed that the plaintext beamforming feedback specified in IEEE 802.11ac/ax—originally designed solely for link adaptation—exposes environment- and device-specific signatures that enable user identification, gesture recognition, and fine-grained localization. Their findings highlight a critical vulnerability: the right-singular subspace of the MIMO channel, when reported verbatim, effectively acts as a stable spatial fingerprint of the surrounding environment.
@@
A recent effort to obfuscate beamforming feedback via randomized angle perturbations [cominelli2024physical] represents an important step forward, but is limited to simple MIMO configurations and lacks a principled privacy interpretation. In particular, the interaction between link adaptation dynamics and injected randomness makes it difficult to isolate and quantify the resulting privacy gains. This motivates the key question we address: Is it possible to design a software-level, standards-compliant CSI feedback mechanism that provides formal privacy guarantees while preserving the beamforming utility of existing Wi-Fi systems?
@@
Summary of Contributions. In this work, we develop a unified framework for privacy-preserving CSI feedback under differential privacy (DP) [dwork2014algorithmic] tailored to the Givens rotation and phase parameters used in IEEE 802.11 compressed beamforming feedback. Instead of modifying the CSI subspace via hidden rotations, we directly privatize the standard-reported angular parameters by replacing deterministic quantization with an $\varepsilon$-DP stochastic quantization rule. The proposed mechanism is fully compatible with the 802.11 feedback structure and admits closed-form sensitivity characterizations for the angular mapping. To the best of our knowledge, this is the first work to
@@
Figure 2: Closed-loop CSI feedback architecture with $N_{t}$ transmit antennas, $N_{r}$ receive antennas, $N_{s}$ spatial streams, a shared angles codebook, and the proposed DP–SQ feedback quantization.
@@ II-C Privacy Threat Model
An Example for Adversarial Speed Estimation. To illustrate the privacy risk, consider the simplest explicit-beamforming configuration with $N_{t}=2$ transmit antennas and $N_{r}=1$ receive antenna (Table. I provides more details on the experimental setup). In this case the right-singular vector of the channel is parameterized by a single Givens rotation angle $\psi_{1,2}$ and two per-antenna phases. The AP uses the quantized angle $\widehat{\psi}_{1,2}(t)$ reported by the STA, but the same value is also observable to a passive eavesdropper. A passive adversary collects a time series of Givens angles $\{\widehat{\psi}_{1,2}(t_{k})\}_{k=1}^{M}$ from successive CSI feedback packets. Mirroring classical micro-Doppler processing used in CSI-based sensing schemes (e.g., [cominelli2024physical]), we now describe how an observer can convert a time series of CSI reports into Doppler and speed estimates. We consider a block-fading MIMO channel with $N_{\mathrm{sc}}$ active subcarriers and $N_{\mathrm{snap}}$ CSI snapshots. At snapshot index $n\in\{0,\dots,N_{\mathrm{snap}}-1\}$, the discrete-time downlink channel is denoted as $\mathbf{H}[n]$, and the corresponding CSI report is generated at time $t_{n}=nT_{\mathrm{CSI}}$, where $T_{\mathrm{CSI}}$ is the CSI reporting interval. In practice, both the access point and a passive eavesdropper operate on an effective CSI obtained by projecting $\mathbf{H}[n]$ onto fixed transmit/receive beams and, in the explicit-beamforming case, reconstructing those beams from compressed CSI feedback. We denote the resulting frequency-domain CSI per subcarrier by
@@
Crucially, this estimation pipeline is agnostic to how the effective CSI $h[k,n]$ is obtained: it can be computed from raw downlink pilots, from beamformed CSI at the receiver, or reconstructed from compressed CSI feedback (e.g., Givens-parameter reports) using the standard IEEE 802.11 beamforming structure. As a result, by continuously tracking the complex CSI associated with the reported beams, a passive eavesdropper can infer the user’s motion speed over time without requiring explicit access the access point’s or STA information as shown Fig. 4 illustrates this attack.
@@ III Stochastic Quantization for Privacy-Preserving Beamforming
In this section, we describe the proposed privacy-preserving mechanism for the beamforming transmit matrix $\mathbf{V}$. The key idea is to apply a randomized quantization procedure  [^1] We want to shed the light that while the receiver noise inherently introduces randomness into the feedback process, it constitutes an uncontrolled and unpredictable source of privacy. Although such noise may provide incidental obfuscation of the beamforming feedback, its distribution and temporal correlation are determined by the physical channel conditions and hardware impairments, and therefore cannot be relied upon to provide formal or quantifiable privacy guarantees. to the Givens–rotation angles that parametrize $\mathbf{V}$ in the IEEE 802.11-style compressed beamforming feedback. Stochastic quantization introduces controlled randomness into each angle, which both masks fine-grained channel structure and enables DP guarantees when the quantization probabilities are appropriately biased. We first review the basic stochastic quantization model and its MSE properties, which will serve as the building block for our proposed DP mechanism.
@@ III-B Differential Privacy
We first assume that the adversary’s observable is the standard feedback payload11footnotetext: While channel sounding may leak some activity-related information, analog CSI is known to weaken significantly with wall penetration and distance, whereas digital beamforming feedback remains robust and observable via passive packet sniffing. : the vector of quantized Givens parameters (phase and rotation angles, or equivalently their codeword indices) derived from the estimated channel. As illustrated in the example in Subsection II-C, an eavesdropper can use these reported quantities to infer user activity or motion patterns directly from the underlying channel realizations.
Definition 1 $($$\varepsilon,\delta$)-DP [dwork2014algorithmic]).
A privacy mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-DP if for any two precoding matrices $\mathbf{V},\mathbf{V}^{\prime}$ and all measurable events $E$ in the output space, the following holds
@@
where $(\epsilon,\delta)$ is the privacy budget such that smaller values of $\epsilon$ and $\delta$ indicate higher levels of privacy and vice versa. The setting when $\delta=0$ is referred as $\varepsilon$-DP.
In the following, we apply this stochastic quantization model to the Givens angles of the beamforming matrix $\mathbf{V}$ and introduce an $\varepsilon$-DP variant by selecting the probability $p$ according to a DP-biased rule.
III-C DP Stochastic Quantization (DP–SQ) Mechanism
Consider a true Givens angle $a$ to be quantized on a uniform grid $\mathcal{Q}$ with spacing $\Delta$, where
@@ III-C DP Stochastic Quantization (DP–SQ) Mechanism
Let $(q_{i},q_{i+1})$ denote the two nearest reconstruction levels to $a$, with circular distance for $\phi$ and clamped distance for $\psi$. The privacy mechanism draws the released value $\widehat{a}$ as
@@
where $\varepsilon_{a}$ is the local privacy budget (e.g., $\varepsilon_{\phi}$ for the phase and $\varepsilon_{\psi}$ for the mixing angle). This randomization satisfies $\varepsilon_{a}$-DP, for any two angles $a,a^{\prime}$. We next present an intermediate lemma to quantify the MSE angular distortion due the DP-SQ mechanism.
@@
The proof is omitted due space limitations. It is worth highlighting that the above expression recovers the MSE guarantee of the deterministic quantization when there is no privacy constraints, i.e., $\varepsilon\rightarrow\infty$.
We next quantify how the DP–SQ scheme of Givens angles perturbs the $N_{s}$-dimensional beamforming subspace. Let $\mathbf{V}^{\star}\!\in\!\mathbb{C}^{N_{t}\times N_{s}}$ denote the ideal right-singular matrix and $\mathbf{P}^{\star}=\mathbf{V}^{\star}(\mathbf{V}^{\star})^{H}$ its associated projector. The quantized precoder generated from DP–SQ angles is denoted $\widehat{\mathbf{V}}$ with projector $\widehat{\mathbf{P}}$. With the DP quantization mechanism established, we now formalize its impact on beamforming performance. The following theorem characterizes the transmit-side subspace distortion induced by our privacy-preserving angular quantization scheme.
Theorem 1 (Expected Subspace Distortion under DP–SQ).
@@
Then the expected squared chordal distance between original subspace $\mathcal{S}^{\star}$ and the perturbed subspace $\widehat{\mathcal{S}}$ via DP–SQ mechanism satisfies:
@@
Using the fact $\sin^{2}x\leq x^{2}$ and the DP–SQ per‐angle MSEs, $\mathbb{E}[\Delta\psi_{i,m}^{2}]=\sigma_{\psi}^{2}$, $\mathbb{E}[\Delta\phi_{i,m}^{2}]=\sigma_{\phi}^{2}$, we obtain

### v2

# Controlling Activity Leakage in Wi-Fi CSI Feedback: A Systematic Design Approach

@@ Abstract
Explicit channel state information (CSI) feedback in IEEE 802.11 conveys transmit beamforming directions by reporting quantized Givens rotation and phase angles that parametrize the right-singular subspace of the channel matrix. Because these angles encode fine-grained spatial signatures of the propagation environment, recent work have shown that plaintext CSI feedback can inadvertently reveal user activity, identity, and location to passive eavesdroppers. In this work, we introduce a standards-compatible differentially private (DP) quantization mechanism that replaces deterministic angular quantization with a stochastic quantizer with a localized differential privacy guarantee, applied directly to the Givens parameters of the transmit beamforming matrix. The mechanism preserves the 802.11 feedback structure, admits closed-form sensitivity bounds for the angular representation, and enables principled privacy calibration. Numerical simulations demonstrate strong privacy guarantees with minimal degradation in beamforming performance.
@@ I Introduction
Beamforming is a core capability of modern multi-antenna wireless systems [6, 5], enabling transmitters to dynamically shape and steer radiated energy toward intended receivers. By coherently combining signals across multiple antennas, the transmitter forms highly directional beams that substantially improve the received signal-to-noise ratio (SNR) while suppressing unintended interference—capabilities unattainable with omnidirectional transmission. Introduced into Wi-Fi through the IEEE 802.11n “High Throughput” amendment [9], beamforming has since become standard in commodity WLAN devices, where the access point (beamformer) computes a channel-dependent steering matrix that optimizes reception quality at the client (beamformee).
Two primary approaches are used to compute this steering matrix [13]. In implicit beamforming, the transmitter estimates the downlink channel from uplink pilot transmissions under an assumption of channel reciprocity. While attractive for its simplicity, this approach is often suboptimal in practice due to hardware asymmetries between the transmit and receive chains. By contrast, explicit beamforming, now the mandated method in modern Wi-Fi standards, achieves higher accuracy by relying on channel state information (CSI) fed back by the client. The beamformee estimates the downlink channel, compresses it, and reports the result to the beamformer through dedicated sounding and feedback frames, typically using quantized matrices or precoding matrix indicators (PMIs).
@@ I-A Privacy Concerns in Explicit Beamforming
The privacy risks associated with channel measurements have been extensively demonstrated in the literature on device-free sensing and Wi-Fi-based activity recognition [1, 15, 10, 14, 16]. These works show that subtle CSI variations can reveal human presence, gestures, and even respiration patterns. However, most existing countermeasures, such as temporal smoothing, random antenna selection, or coarse quantization, offer only heuristic protection, lack formal privacy guarantees, and often degrade beamforming performance.
More recently, empirical studies have revealed that standard Wi-Fi beamforming feedback itself can directly leak sensitive information. Liu et al. [11] showed that the plaintext beamforming feedback specified in IEEE 802.11ac/ax, originally designed solely for link adaptationm exposes environment- and device-specific signatures that enable user identification, gesture recognition, and fine-grained localization. Their findings highlight a critical vulnerability: the right-singular subspace of the MIMO channel, when reported verbatim, effectively acts as a stable spatial fingerprint of the surrounding environment.
I-B Related Work
@@
A recent effort to obfuscate beamforming feedback via randomized angle perturbations [2] represents an important step forward, but is limited to simple MIMO configurations and lacks a principled privacy interpretation. In particular, the interaction between link adaptation dynamics and injected randomness makes it difficult to isolate and quantify the resulting privacy gains. This motivates the key question we address: Is it possible to design a software-level, standards-compliant CSI feedback mechanism that provides formal privacy guarantees while preserving the beamforming utility of existing Wi-Fi systems?
@@
Summary of Contributions. In this work, we develop a unified framework for privacy-preserving CSI feedback under differential privacy (DP) [3] tailored to the Givens rotation and phase parameters used in IEEE 802.11 compressed beamforming feedback. Instead of modifying the CSI subspace via hidden rotations, we directly privatize the standard-reported angular parameters by replacing deterministic quantization with an $\varepsilon$-DP stochastic quantization rule. The proposed mechanism is fully compatible with the 802.11 feedback structure and admits closed-form sensitivity characterizations for the angular mapping. To the best of our knowledge, this is the first work to
@@
Figure 2: Closed-loop CSI feedback architecture with $N_{t}$ transmit antennas, $N_{r}$ receive antennas, $N_{s}$ spatial streams, a shared angles codebook, and the proposed DP-SQ feedback quantization.
@@ II-C Privacy Threat Model
An Example for Adversarial Speed Estimation. To illustrate the privacy risk, consider the simplest explicit-beamforming configuration with $N_{t}=2$ transmit antennas and $N_{r}=1$ receive antenna (Table. I provides more details on the experimental setup). In this case the right-singular vector of the channel is parameterized by a single Givens rotation angle $\psi_{1,2}$ and two per-antenna phases. The AP uses the quantized angle $\widehat{\psi}_{1,2}(t)$ reported by the STA, but the same value is also observable to a passive eavesdropper. A passive adversary collects a time series of Givens angles $\{\widehat{\psi}_{1,2}(t_{k})\}_{k=1}^{M}$ from successive CSI feedback packets. Mirroring classical micro-Doppler processing used in CSI-based sensing schemes (e.g., [2]), we now describe how an observer can convert a time series of CSI reports into Doppler and speed estimates. We consider a block-fading MIMO channel with $N_{\mathrm{sc}}$ active subcarriers and $N_{\mathrm{snap}}$ CSI snapshots. At snapshot index $n\in\{0,\dots,N_{\mathrm{snap}}-1\}$, the discrete-time downlink channel is denoted as $\mathbf{H}[n]$, and the corresponding CSI report is generated at time $t_{n}=nT_{\mathrm{CSI}}$, where $T_{\mathrm{CSI}}$ is the CSI reporting interval. In practice, both the access point and a passive eavesdropper operate on an effective CSI obtained by projecting $\mathbf{H}[n]$ onto fixed transmit/receive beams and, in the explicit-beamforming case, reconstructing those beams from compressed CSI feedback. We denote the resulting frequency-domain CSI per subcarrier by
@@
Crucially, this estimation pipeline is agnostic to how the effective CSI $h[k,n]$ is obtained: it can be computed from raw downlink pilots, from beamformed CSI at the receiver, or reconstructed from compressed CSI feedback (e.g., Givens-parameter reports) using the standard IEEE 802.11 beamforming structure. As a result, by continuously tracking the complex CSI associated with the reported beams, a passive eavesdropper can infer the user’s motion speed over time without requiring explicit access the access point’s or STA information as shown Fig. 4 illustrates this attack. Because our approach operates exclusively on the standard Givens rotation and phase angles already utilized in IEEE 802.11 explicit feedback, it requires no hardware modifications or alterations to existing channel sounding procedures. Consequently, it serves as a seamless, drop-in privacy enhancement for commercial transceivers that rigorously preserves the required orthonormality and structure of the beamforming matrix.
@@ III Stochastic Quantization for Privacy-Preserving Beamforming
In this section, we describe the proposed privacy-preserving mechanism for the beamforming transmit matrix $\mathbf{V}$. The key idea is to apply a randomized quantization procedure  [^1] We want to shed the light that while the receiver noise inherently introduces randomness into the feedback process, it constitutes an uncontrolled and unpredictable source of privacy. Although such noise may provide incidental obfuscation of the beamforming feedback, its distribution and temporal correlation are determined by the physical channel conditions and hardware impairments, and therefore cannot be relied upon to provide formal or quantifiable privacy guarantees. to the Givens–rotation angles that parametrize $\mathbf{V}$ in the IEEE 802.11-style compressed beamforming feedback. Stochastic quantization introduces controlled randomness into each angle, which both masks fine-grained channel structure and enables formal DP guarantees when the quantization probabilities are appropriately biased. We first review the basic stochastic quantization model and its MSE properties, which will serve as the building block for our proposed DP mechanism.
@@ III-B Differential Privacy
We consider an adversary whose observable is the standard feedback payload11footnotetext: While channel sounding may leak some activity-related information, analog CSI is known to attenuate significantly with wall penetration and distance, whereas digital beamforming feedback remains robust and can be passively observed via packet sniffing. : namely, the vector of quantized Givens parameters (phase and rotation angles, or equivalently their codeword indices) derived from the estimated channel. As illustrated in Subsection II-C, an eavesdropper can exploit these reported quantities to infer user activity or motion patterns from the underlying channel realizations.
We next recall the standard notion of differential privacy (DP), which provides a formal measure of indistinguishability between outputs corresponding to neighboring inputs.
Definition 1 $($$\varepsilon,\delta$)-DP [3]).
A randomized mechanism $\mathcal{M}$ is said to be $(\varepsilon,\delta)$-differentially private if for any two inputs $\mathbf{V},\mathbf{V}^{\prime}$ and any measurable set $E$ in the output space,
@@
where $\varepsilon>0$ and $\delta\geq 0$ are privacy parameters. Smaller values of $(\varepsilon,\delta)$ correspond to stronger privacy guarantees. The special case $\delta=0$ is referred to as pure $\varepsilon$-DP.
In the context of beamforming feedback, directly enforcing global $(\varepsilon,\delta)$-DP over the continuous angular domain would require full-support randomization across all reconstruction levels, which can significantly degrade the utility of the reported beamforming parameters. Instead, we adopt a localized privacy notion aligned with the quantized representation of the feedback, as detailed next.
III-C DP Stochastic Quantization (DP-SQ) Mechanism
Consider a true Givens angle $a$ to be quantized on a uniform grid $\mathcal{Q}$ with spacing $\Delta$. For the phase and mixing angles, we have
@@ III-C DP Stochastic Quantization (DP-SQ) Mechanism
Let $(q_{i},q_{i+1})$ denote the two nearest reconstruction levels to $a$, i.e.,
$q_{i}\leq a\leq q_{i+1},\quad i\in\{1,\dots,|\mathcal{Q}|-1\},$
where distances are interpreted in the circular sense for $\phi$ and in the standard (clamped) sense for $\psi$.
We define a local adjacency relation on the angular domain: two angles $a$ and $a^{\prime}$ are said to be adjacent, denoted $a\sim a^{\prime}$, if they belong to the same quantization interval, $a,a^{\prime}\in[q_{i},q_{i+1}]\quad\text{for some }i$.
[…]
@@
where $\varepsilon_{a}$ is a design parameter controlling the level of randomization (e.g., $\varepsilon_{\phi}$ for the phase and $\varepsilon_{\psi}$ for the mixing angle).
The above mechanism preserves the standard two-level stochastic quantization structure while introducing controlled randomization between adjacent reconstruction levels. In particular, the released value $\widehat{a}$ depends only on the local quantization cell $q_{i},q_{i+1}]$ and not on the exact location of $a$ within the cell.
Privacy guarantee. Under the cell-wise adjacency relation, the DP-SQ mechanism provides the following privacy guarantee.
The above guarantee implies that all angles within the same quantization interval are statistically indistinguishable, thereby limiting the resolution at which the underlying angle can be inferred. The parameter $\varepsilon_{a}$ governs a local privacy–utility tradeoff by controlling the bias between the two adjacent reconstruction levels: as $\varepsilon_{a}\to 0$, the mechanism approaches uniform randomization, while as $\varepsilon_{a}\to\infty$, it approaches deterministic quantization. We emphasize that this privacy notion is local to each quantization cell. Achieving global $\varepsilon$-local DP over arbitrary input pairs would require full-support randomization across all reconstruction levels (e.g., via $M$-ary randomized response), which would significantly increase angular distortion and degrade beamforming performance. The proposed DP-SQ mechanism instead focuses on localized randomization that preserves the structure and efficiency of standard CSI feedback representations.
Utility. We next introduce an intermediate lemma to quantify the distortion introduced by the DP-SQ mechanism.
@@
It is worth highlighting that the above expression recovers the MSE guarantee of the deterministic quantization when there is no privacy constraints, i.e., $\varepsilon\rightarrow\infty$.
Proof.
Consider a uniform mid-rise quantizer with step size $\Delta$. For a given true angle $a$, let $r$ denote its distance to the nearest quantization level. Then $r:=\operatorname{dist}(a,\text{nearest level})\in[0,\Delta/2]$. Under the standard high–resolution assumption, $a$ is uniformly distributed within its quantization cell, so $r\sim\mathrm{Unif}\bigl([0,\Delta/2]\bigr)$. With privacy parameter $\varepsilon>0$, the DP stochastic quantizer outputs the nearest level with probability $p^{\star}(\varepsilon)$, and the farther neighbor (at distance $\Delta-r$) with probability $1-p^{\star}(\varepsilon)$. Conditioned on $r$, the MSE is
$\mathbb{E}\big[(\widehat{a}-a)^{2}\mid r\big]=p^{\star}(\varepsilon)\,r^{2}+\bigl(1-p^{\star}(\varepsilon)\bigr)(\Delta-r)^{2}.$
The pdf of $r$ is $f_{r}(x)=\tfrac{2}{\Delta}$ on $0,\Delta/2]$, hence
$\mathbb{E}[r]$ $=\int_{0}^{\Delta/2}x\,\cdot\frac{2}{\Delta}\,dx=\frac{2}{\Delta}\cdot\frac{x^{2}}{2}\Big|_{0}^{\Delta/2}=\frac{\Delta}{4},$
$\mathbb{E}[r^{2}]$ $=\int_{0}^{\Delta/2}x^{2}\,\cdot\frac{2}{\Delta}\,dx=\frac{2}{\Delta}\cdot\frac{x^{3}}{3}\Big|_{0}^{\Delta/2}=\frac{\Delta^{2}}{12}.$
Similarly,
$\mathbb{E}\big[(\Delta-r)^{2}\big]$ $=\Delta^{2}-2\Delta\,\mathbb{E}[r]+\mathbb{E}[r^{2}]$
$=\Delta^{2}-\frac{\Delta^{2}}{2}+\frac{\Delta^{2}}{12}=\frac{7\Delta^{2}}{12}.$
Now average $\mathbb{E}[(\widehat{a}-a)^{2}\mid r]$ over $r$, we get
$\mathbb{E}\big[(\widehat{a}-a)^{2}\big]$ $=p^{\star}(\varepsilon)\,\mathbb{E}[r^{2}]+\bigl(1-p^{\star}(\varepsilon)\bigr)\,\mathbb{E}\big[(\Delta-r)^{2}\big]$
$=\frac{\Delta^{2}}{12}\,\cdot\bigl(7-6\,p^{\star}(\varepsilon)\bigr).$
Substituting $p^{\star}(\varepsilon)=\tfrac{1+\kappa(\varepsilon)}{2}$ gives
$7-6\,p^{\star}(\varepsilon)$ $=7-3\bigl(1+\kappa(\varepsilon)\bigr)=4-3\,\kappa(\varepsilon),$
so
$\mathbb{E}\big[(\widehat{a}-a)^{2}\big]=\frac{\Delta^{2}}{12}\,\cdot\bigl(4-3\,\kappa(\varepsilon)\bigr),$
which is the claimed expression. ∎
Fig. 5 illustrates the variability induced by the proposed DP-SQ mechanism relative to deterministic quantization. In contrast to the deterministic scheme, which yields a fixed PMF over reconstruction levels, DP-SQ introduces stochasticity through randomized selection between adjacent quantization points. The resulting mean PMF and its associated standard deviation show that DP-SQ preserves the dominant structure of the deterministic angle distribution while causing only a localized spreading of probability mass around nearby reconstruction levels.
Figure 5: Empirical PMFs of the quantized Givens angles under deterministic quantization and the proposed DP-SQ mechanism for single-stream beamforming with $(N_{t},N_{r})=(4,2)$, $(B_{\phi},B_{\psi})=(6,4)$, privacy parameters $(\varepsilon_{\phi},\varepsilon_{\psi})=(0.8,0.8)$, and $T=100$ independent DP-SQ realizations. Each subplot corresponds to one reported angle in the compressed feedback parameterization. The gray bars denote the deterministic PMF over reconstruction levels, while the red curves represent the mean PMF produced by DP-SQ across repeated runs, and the shaded regions indicate one standard deviation. The underlying raw angle samples are fixed across runs, and only the DP-SQ mechanism is randomized, thereby isolating the variability due solely to privacy-induced randomness.
We next quantify how the DP-SQ scheme of Givens angles perturbs the $N_{s}$-dimensional beamforming subspace. Let $\mathbf{V}^{\star}\!\in\!\mathbb{C}^{N_{t}\times N_{s}}$ denote the ideal right-singular matrix and $\mathbf{P}^{\star}=\mathbf{V}^{\star}(\mathbf{V}^{\star})^{H}$ its associated projector. The quantized precoder generated from DP-SQ angles is denoted $\widehat{\mathbf{V}}$ with projector $\widehat{\mathbf{P}}$. With the DP quantization mechanism established, we now formalize its impact on beamforming performance. The following theorem characterizes the transmit-side subspace distortion induced by our privacy-preserving angular quantization scheme.
Theorem 1 (Expected Subspace Distortion under DP-SQ).
@@
Then the expected squared chordal distance between original subspace $\mathcal{S}^{\star}$ and the perturbed subspace $\widehat{\mathcal{S}}$ via DP-SQ mechanism satisfies:
@@
Using the fact $\sin^{2}x\leq x^{2}$ and the DP-SQ per‐angle MSEs, $\mathbb{E}[\Delta\psi_{i,m}^{2}]=\sigma_{\psi}^{2}$, $\mathbb{E}[\Delta\phi_{i,m}^{2}]=\sigma_{\phi}^{2}$, we obtain